# An Online Data Analysis Framework for Accelerator-Based Physics Experiments

### v2

@@ 3.1.1 Challenges with Timestamp Tagging
Initially, the process of sorting devices’ data via timestamp was to be carried out by automatically generating duplicates of the devices’ output data files and appending a UNIX timestamp [37] to the filename; this proposed (but ultimately unsuccessful) process is outlined in Figure 1. The timestamp used to label device data was to be provided by CERN’s Network Time Protocol (NTP) [38], and corresponds to the timestamp used to index CERN’s accelerator cycles. This timestamp was termed the “cyclestamp”. Ultimately, the framework’s code (see Section 4) was built to select shots by the UNIX timestamps provided by the secondary log, rather than by the CERN NTP cyclestamp.
A revised scheme (shown in Figure 2) was introduced to compensate for the difficulties that arose while attempting to implement the scheme shown in Figure 1 due to unreliability of the remote connection between local devices running the framework, and servers using the CERN NTP. In place of an automated log of timestamps at which each device triggered, a simplified “secondary shot log” was maintained, which indexed the data from each device by the UNIX timestamp corresponding to the upload time on CERNbox.
@@ 4 Code Structure and Design
The goal of this project was to construct a set of analysis codes which could perform analysis on the output data for all devices on the experiment as soon as the data was available on a centralized repository. This demanded both flexibility (the ability to introduce and remove diagnostics at will) and specificity (with each device requiring its own unique analysis and processing). The codebase was written in Python 3.11.9.
The code’s hierarchical structure, shown schematically in Figure 3, was chosen to allow a high degree of modularity, addressing both the issue of flexibility and the need to specify analysis individually for each device.
The code’s high degree of modularity and class-based structure also allowed members of the collaboration not directly involved in the code’s development to sample code piecewise from the analysis pipeline to perform their own specialized analysis in parallel with the online data collection.
@@ 4.1 Configurability
The code was designed to read commands from an input configuration (either via JSON file or a manually-edited “input configuration” dictionary in the Python environment) to maximize readability and customization. The JSON format is simple, has minimal syntax, and closely resembles plain English. The configuration file was its ability to reduce the proliferation of hard-coded values in the framework which would need to be changed between runs, a process prone to human error. In the input configuration, users could specify shot numbers of interest and/or request statistical averaging across these shots. As depicted in Figure 4, the user is able to specify “background” data to be subtracted from data of interest. The configuration allowed for displaying images without performing any analysis to expedite execution speed; quantitative assessment of how this contributed to the code’s execution speed can be found in table 2.
@@ 4.2 Modularity
The code’s modular structure, depicted in Figure 3, was key to its adaptability. As experimental goals were modified, or regions of interest identified, quick refactors could be made to analysis methods without disrupting the data collection and extraction processes. This allowed for a “plug-in-and-play” model: collaborators external to the project could implement their own analysis code, done independently and without any knowledge of the HRMT-68 online analysis framework, by simply inserting their extraction and analysis methods into the relevant modules in the framework.
The file manager module is tasked with creating an index of file names and their corresponding upload time, and this index is converted into a time-ordered index of file paths by the path manager. The run and load manager modules have two subclasses, one each for handling time-domain waveform data from oscilloscopes, and for handling bitmap image data from cameras. All device-specific analysis during runtime is controlled by the run manager, including statistical averaging and error propagation. The load manager subclass that the code calls is device-specific to account for the different file types and data logging structure of each device. All analysis and data visualization after the relevant data is extracted and preprocessed is handled by the operations manager, after which the code terminates.
@@ 5 Performance and Results
In this section, we quantify the code’s performance when producing analysis and/or visualization of diagnostic device data. For these tests, we tested the code with data from: the OTR streak unit camera, synchrotron spectrometer camera, a single chromox camera (the plasma cell camera monitoring the upstream chromox screen, see Section 2), and oscilloscope 1 (see Section 1).
In Section 5.1, we present a summary of the code’s performance in Tables 2 and 3, which measure execution time and peak memory usage, respectively. The tests were performed over a remote connection to a CPU (Intel(R) Xeon(R) Silver 4216 CPU, 2.10 GHz [39]) on the CERN cluster, running an AlmaLinux9 OS distribution [40], via CERN’s SWAN service, with a cumulative 8 GB of RAM and 2 cores allocated to the session. The ethernet connection of the remote CPU was configured to a ceiling of 10000 Mbps, and measurement of computer’s download speed yielded . Remote connection to the remote cluster CPU via the SWAN service was achieved over a Wi-Fi connection.
All tests were performed with the background subtraction routine - this was done by averaging over 5 shots designated as background to create a composite background image, which was then subtracted from the foreground image. This is in order to demonstrate the speed of execution even when a larger computational load is demanded of the code. For shot averaging, background removal was also performed, with the same 5 background shots selected to produce the background composite. The foreground image was produced by averaging over five shots, meaning that the code was required to process 10 images during runs with foreground averaging. Each entry of execution time in Table 2 is an average over 10 trials.
As mentioned previously in section 4.1, the code could be configured to perform no analysis and simply plot the image, in order to increase execution speed in cases where it was only necessarily to visualize the camera images or oscilloscope traces. Therefore, tests of execution speed were also extended to cases both with and without analysis functions being called by the code. Results for tests of peak memory usage in Table 3 were averaged across 10 runs of the code. This was measured using Python 3’s inbuilt memory profiler/memory usage function.
In order to account for the significant reduction in the code’s speed introduced by calling analysis methods on the chromox cameras, additional tests to quantify two of the most computationally-expensive analysis methods, those related to calculating first- and second-moments and performing coordinate transformations (Section 2.2), involved in the chromox calculations were performed. These are summarized in Table 4.
@@ 5.1 Results and Discussion
From the tests of code speed displayed in Table 2, we conclude that the foreground-averaging in the absence of analysis methods introduces negligible additional overhead for the chromox camera and B-dot probe data. In the former case, execution time for foreground-averaging only introduces a increase in execution time, whereas effects to execution time for displays on the B-dot probe oscilloscope execution time are statistically negligible, and indeed execution time actually decreased, changing by when averaging was introduced. We conclude that the code’s statistical averaging methods are not the limiting factor, and that instead variations in connection/download speed between clusters accessed via SWAN and the CERNbox service, affected by network load and user demand, are far more likely to be responsible. However, due to user restrictions on the SWAN service, network round-trip time (RTT) between SWAN and CERNbox could not be tested via pinging the SWAN domain.
Inclusion of analysis methods, however, did introduce significant overhead in the execution process- in the case of the chromox cameras: this resulted in percentage increases of and in execution time for the “single shot” and “average shot” cases, respectively. The methods used to calculate first- and second-order moments are computationally-costly (see Section 2.2), iterating over all pixels in the image and therefore scaling as for an image of pixel dimension . Additionally, in order to display radial/azimuthal marginals, methods tasked with transforming the images’ pixels from Cartesian coordinates, , to planar polar coordinates, , were implemented. As shown in Table 4, the execution times for these two functions were investigated individually. Cumulatively, they accounted for of the entire runtime for calling analysis on the chromox images.
@@
Operations with the streak unit camera were clear outliers in both execution time (Table 2) and memory consumption (Table 3). This is likely due to the larger pixel dimensions of the images from the synchrotron spectrometer and OTR streak unit cameras ( and pixels, respectively) compared to the dimensions of the single chromox camera which was used during code testing ( pixels).
@@
For streak unit analysis, variation between cases with and without analysis was larger ( variation for the “single shot” case (once again, a decrease in execution time), and increase for the “average shot” case). It is worth noting that the comparatively large variation between cases with and without analysis for the “single shot” case is unlikely representative of inefficiencies in the code; execution time was shown to actually decrease when analysis was introduced, despite the fact that operations performed by the code without calling analysis are a subset of operations performed when analysis is requested.
We also highlight that, for all devices, the analysis methods incur only a small amount of memory cost; indeed, memory consumption during tests across all devices actually decreased between the “singleshot” cases when analysis was introduced; owing to the fact that there was an increase in memory consumption after the addition of analysis for the “averageshot” cases, and the fact that the drop in peak memory consumption occurred for all devices, we suspect that this may be due to changes in the Jupyter kernel state between tests, though the exact mechanism explaining this occurrence remains unknown. This may also explain the decrease in execution time when analysis was introduced in the “singleshot” cases for the OTR streak unit camera and spectrometer cameras, which are difficult to analyze due to the increased number of operations performed in the analysis case.
@@ 6 Conclusions and Future Work
In this work, we have presented an online framework which is flexible to varying demands from a user, is moderately fast even under comparatively adversarial circumstances (connection to a remote server over a wireless connection with modest download speed), and is highly modular, allowing easy implementation of data extraction and analysis code for additional diagnostic devices. Though the code’s performance during experimental runtime was largely successful, we conclude by highlighting the framework’s weaknesses, and how the framework could be improved and modified.
@@ 6.1 Introducing Caching
Improvements could be made to execution speed by locally caching data from shots of interest. Often, it was desirable to plot data from the same shot multiple times in order to, for example, crop to regions of interest or adjust image contrast for visibility. This required multiple code runs, with each run limited by the speed at which data could be acquired over a network connection. In future, it would be useful to introduce an optional “caching” feature, allowing users to store data from a desired shot locally over a short period, with the option of “clearing” the cache when this data was no longer required. Caching was performed in only one case, and this was for combining data from the two different oscilloscopes used during runtime in order to allow subtraction between the two channels of the downstream longitudinal B-dot probe (see Section 2.1).
@@ 6.2 Tagging by Cyclestamp
Experimental collaborations using the framework should aim to have some universally-recognized system of logging data by acquisition time. At CERN, as discussed in Section 3, a strong candidate would be the UNIX timestamp corresponding to each accelerator cycle, and data filenames can be “stamped” by unique UNIX number. The framework codebase is easily adjusted to allow indexing by UNIX timestamp so long as it knows the position of the timestamp in the filename string; however, due to problems highlighted in Section 3, this was not implemented.
@@ 6.3 Avoiding Intermediate Cloud-Based Services
Though the CERN SWAN service allowed direct user access to the CERNbox service, it introduced several obstacles. SWAN has no native functionality for editing .json file types, and thus configuration had to be stored manually in a front-facing Jupyter notebook file. The Jupyter kernel introduces further overhead, and online code refactors in the framework modules are not recognized unless the kernel is restarted. Additionally, SWAN requires a stable connection to the CERN network, which is in-demand during working hours and this introduces unpredictable latency.
Future experiments should aim to save data to a local computer on which the framework can be used, or at least maintain a frequently-updated clone of the data repository for ease of access which does not require an internet connection during code execution.

### v3

@@ 3.1.1 Challenges with Timestamp Tagging
A central goal of the framework was the creation of a consistent scheme for logging all data acquired during the experiment. During the HRMT-68 experiment, several experimental conditions were continually modified. These included power supplied to the plasma cell, position of the graphite target used to produce the electron-positron beam, as well as the beam density requested from the SPS. Keeping track of which data files corresponded to which experimental conditions was therefore essential in subsequent data analysis. Therefore, we aimed to maintain a centralized “shot log” which tracked all experimental conditions during each beam extraction, and the time of each extraction. The aim was to then compare these extraction times to the time(s) at which each device was triggered.
Initially, the process of sorting devices’ data via timestamp was to be carried out by automatically generating duplicates of the devices’ output data files and appending a UNIX timestamp [louis2020time] to the filename. This proposed process is outlined in Figure 1. The timestamp used to label device data was to be provided by CERN’s Network Time Protocol (NTP) [K2023], and corresponds to the timestamp used to index CERN’s accelerator cycles. This timestamp is termed the “cyclestamp”. During our data collection and logging procedures, when relying on a connection the CERN NTP server, we observed up to an discrepancy between the timestamps associated with each devices’ data for the same extraction. This limitation was due to delays in the devices’ data being saved to our cloud-based repository. However, this observed precision was more than sufficient to remove ambiguity surrounding which data corresponded to which extraction, as intervals between subsequent extractions from the beamline during the experimental run were around 90 seconds.
However, this initial method of tagging the filename with a timestamp directly proved infeasible for several reasons: firstly, it necessitates duplication of all data, resulting in a doubling of storage requirements. Given that hard disk space is often at a premium in small-scale experiments, this complication is unsatisfactory and will often be unacceptable. Secondly, timestamping the files (by creating a copies of each file with a modified filename) relied on continuous execution of a Python script which maintains a connection to the CERN NTP server via a WebSocket connection [rfc6455], which increased the framework’s vulnerability to endemic network latency and connectivity concerns. As a result, only a fraction of collected data was correctly timestamped. We retain a description of this idealized process in order to motivate future efforts to dynamically tag data in such an unambiguous fashion while addressing storage and connectivity concerns.
A revised scheme (shown in Figure 2) was introduced to compensate for the difficulties that arose while attempting to implement the scheme shown in Figure 1. In place of an automated log of timestamps at which each device triggered, a simplified “secondary shot log” was maintained, which indexed the data from each device by the UNIX timestamp corresponding to the upload time on CERNbox rather than by relying on the cyclestamp. Unlike the central “master” log, which had to be maintained by hand and contained details of experimental conditions, the secondary log contained no detail of experimental conditions, such as target placement or conditions of the plasma. However, the secondary shot log was automatically generated, and hence resilient to human error.
This process, while cumbersome, was internally consistent and resulted in the creation of two centralized shot-logs for all diagnostic devices, where data files were indexed by a universal timestamp key. This eliminated ambiguity as to which data files corresponded to which experimental conditions and which beam extraction. These timestamp keys can then be fed into the framework’s frontend to filter for data of interest across all devices.
Our framework retains the ability to access data for either method, both the case in which the filename is explicitly modified and appended with the UNIX timestamp, or the case in which data is logged by the time at which it first appeared on the EOS/CERNbox file system. The former case is preferable, as it provides all data with a permanent timestamp which is resilient to file modification or migration of data off the EOS/CERNbox storage system; both cases will result in changes to the files’ modification time metadata, which was used in our revised method (Figure 2) to index data.
@@ 4 Code Structure and Design
Having resolved the issue of indexing data by the time of acquisition, allowing us to filter out shots of interest, the subsequent goal of this project was to construct a set of analysis methods to be performed on the output data across all devices. Our desire to do so “online” meant that these methods had to be executable as soon as the data was available on our centralized CERNBOX/EOS repository. This demanded both flexibility (the ability to introduce and remove diagnostics at will) and specificity (with each device requiring its own unique analysis and processing). The codebase was written in Python 3.11.9.
The choice of Python resulted in somewhat of a compromise on speed and efficiency. However, it provided equally significant gains in flexibility. Python is by some metrics the most widely-used programming language [TIOBE_Index]; hence our framework necessitates no specialist skills other than basic proficiency in a common programming language, compared to the specialist expertise required for device integration into NXCALS/FESA. This guarantees a baseline level of longevity (Python is unlikely to exit use in academic circles in the near future) and accessibility. Moreover, Python offers an extensive array of packages specialized in data extraction and analysis, such as the popular Pandas package [reback2020pandas] and the Numpy library, whose performance is optimized via the use of an underlying C/C++ backend for accelerated computation [harris2020array].
The code’s hierarchical structure, shown schematically in Figure 3, was chosen to allow a high degree of modularity, addressing both the issue of flexibility and the need to specify analysis individually for each device. Its high degree of modularity and class-based structure also allowed members of the collaboration not directly involved in the code’s development to sample code piecewise from the analysis pipeline to perform their own specialized analysis in parallel with the online data collection.
@@ 4.1 Configurability
The code was designed to read commands from an input configuration (either via JSON file or a manually-edited “input configuration” dictionary in the Python environment) to maximize readability and customization. The JSON format is simple, has minimal syntax, and closely resembles plain English. The purpose of the configuration file was to reduce the proliferation of hard-coded values in the framework which would need to be modified between runs, a process prone to human error. In the input configuration, users could specify shot numbers of interest and/or request statistical averaging across these shots. As depicted in Figure 4, the user is able to specify “background” data to be subtracted from data of interest. The configuration allowed for displaying images without performing any analysis to expedite execution speed; quantitative assessment of how this contributed to the code’s execution speed can be found in Table 3.
@@ 4.2 Modularity
The code’s modular structure, depicted in Figure 3, focusing on the image analysis functionality, was key to its adaptability. As experimental goals were modified, or regions of interest identified, quick refactors could be made to analysis methods without disrupting the data collection and extraction processes. This allowed for a “plug-in-and-play” model: collaborators external to the project could implement their own analysis code, done independently and without any knowledge of the HRMT-68 online analysis framework, by simply inserting their extraction and analysis methods into the relevant modules in the framework. A detailed description of the different key-modules are described below.
The file manager module is tasked with creating an index of file names and their corresponding upload time, and this index is converted into a time-ordered index of file paths by the path manager, as described in Section 3. The run and load manager modules have two subclasses, one each for handling time-domain waveform data from oscilloscopes, and for handling bitmap image data from cameras. All device-specific analysis during runtime is controlled by the run manager, including statistical averaging and error propagation. The load manager subclass that the code calls is device-specific to account for the different file types and data logging structure of each device. All analysis and data visualization after the relevant data is extracted and preprocessed is handled by the operations manager, after which the code terminates.
This hierarchical substructure is ideal as it allows for the straightforward, “plug-in” implementation of analysis methods for additional diagnostic devices, as required. As new devices are introduced to an experimental setup, desired analysis methods, such as visualization, cropping, and marginalization over an axis, can be implemented in the operations manager module. This can be done without modification to the path, run, and load manager classes, which contain generic functionality for logging and extracting data from any device.
@@ 5 Performance and Results
In this section, we quantify the code’s performance when producing analysis and/or visualization of diagnostic device data. For these tests, we tested the code with data from: the OTR streak unit camera, synchrotron spectrometer camera, a single Chromox camera (a single plasma cell camera, see Section 2), and an oscilloscope reading output from one of our B-Dot probes (see Section 1).
In Section 5.1, we present a summary of the code’s performance in Tables 3 and 4, where we measure execution time and peak memory usage, respectively. The tests were performed over a remote connection to a CPU (Intel(R) Xeon(R) Silver 4216 CPU, 2.10 GHz [intel_xeon]) on the CERN cluster, running an AlmaLinux9 OS distribution [almalinux], via CERN’s SWAN service, with a cumulative 8 GB of RAM and 2 cores allocated to the session. The ethernet connection of the remote CPU was configured to a ceiling of 10000 Mbps, and measurement of computer’s download speed yielded .
All tests were performed with the background subtraction routine to maximize computational complexity. This was produce estimates for a “worst-case” run time during testing. Background subtraction was done by averaging over 5 shots designated as “background” to create a “composite background” image, which was then subtracted from the foreground image (see Figure 4). Summaries of pixel dimensions for the three types of device producing visual data are provided in Table 2
| Device |
| Synchrotron Spectrometer | |||||||||||
| OTR Streak Camera | |||||||||||
| Chromox Camera |
Each entry of execution time in Table 3 is an average over 10 trials.
As mentioned previously in section 4.1, the code could have its configuration modified between runs. Users could specify that the framework perform no analysis and simply plot device data in order to increase execution speed in cases where it was only necessarily to visualize the camera images or oscilloscope traces, and we include such test cases in our performance evaluation. We provide a sample output from the analysis framework for one of the Chromox cameras (see Section 2), HRM3, in Figure 5.
Results for tests of peak memory usage in Table 4 were averaged across 10 runs of the code. This was measured using Python 3’s inbuilt memory profiler/memory usage function.
In order to account for the significant reduction in the code’s speed introduced by calling analysis methods on the Chromox cameras, additional tests to quantify two of the most computationally-expensive analysis methods involved in the Chromox calculations were performed. These computationally expensive methods included calculating the first and second moments of image pixel intensity (i.e. the centroid of pixel intensity, and standard deviations on this centroid in both dimensions), as well as coordinate transformations from 2D Cartesian coordinates to plane polar coordinates. The latter method was performed in order to generate radial and azimuthal marginals. These are summarized in Table 5.
@@ 5.1 Results and Discussion
From the tests of code speed displayed in Table 3, we conclude that the foreground-averaging in the absence of analysis methods introduces negligible additional overhead for the Chromox camera and B-dot probe data. In the former case, execution time for foreground-averaging only introduces a increase in execution time, whereas effects to execution time for displays on the B-dot probe oscilloscope execution time are statistically negligible, and indeed execution time actually decreased, changing by when averaging was introduced. We conclude that the code’s statistical averaging methods are not the limiting factor, and that instead variations in connection/download speed between clusters accessed via SWAN and the CERNbox service, affected by network load and user demand, are far more likely to be responsible.
Inclusion of analysis methods, however, did introduce significant overhead in the execution process; in the case of the Chromox camera images: this resulted in percentage increases of and in execution time for the “single shot” and “average shot” cases, respectively. The methods used to calculate first- and second-order moments are computationally-costly, iterating over all pixels in the image and therefore scaling as for an image of pixel dimension . Transforming the images’ pixels from Cartesian to planar polar coordinates also shared this computational complexity. As shown in Table 5, the execution times for these two functions were investigated individually. Cumulatively, they accounted for of the entire runtime for calling analysis on the Chromox images.
@@
Operations with the streak unit camera were clear outliers in both execution time (Table 3) and memory consumption (Table 4). This is likely due to the larger pixel dimensions of the images from the synchrotron spectrometer and OTR streak unit cameras ( and pixels per image, respectively) compared to the dimensions of the single Chromox camera which was used during code testing ( pixels).
@@
For streak unit analysis, variation between cases with and without analysis was larger ( variation for the “single shot” case (once again, a decrease in execution time), and increase for the “average shot” case). It is worth noting that the comparatively large variation between cases with and without analysis for the “single shot” case is unlikely representative of inefficiencies in the code; execution time was shown to actually decrease when analysis was introduced, despite the fact that operations performed by the code without calling analysis are a subset of operations performed when analysis is requested. This is most likely due to variations in network conditions.
We would also like to highlight that, for all devices, the analysis methods incur only a small amount of memory cost; indeed, memory consumption during tests across all devices actually decreased between the “single shot” cases when analysis was introduced; owing to the fact that there was an increase in memory consumption after the addition of analysis for the “average shot” cases, and the fact that the drop in peak memory consumption occurred for all devices, we suspect that this may be due to changes in the Jupyter kernel state between tests, though the exact mechanism explaining this occurrence remains unknown. This may also explain the decrease in execution time when analysis was introduced in the “single shot” cases for the OTR streak unit camera and spectrometer cameras, which are difficult to analyze due to the increased number of operations performed in the analysis case.
5.2 Evaluation
Though our framework’s performance is not state-of-the art, its ability to address limitations on data collection (as described in Section 3) and cater to a wide range of diagnostics (Section 2) while maintaining a high level of flexibility and customizability made it invaluable to the experimental runtime. Our decision to implement our analysis methods directly in the framework’s codebase also served as an invaluable, internal “alpha test” phase of our analysis scripts. For example, we were able to identify regions and shots of interest in preparation for our extensive offline data analysis efforts; thus, such specialist frameworks have significant potential in providing small-scale experimental collaborations with a “running start” on their analysis for subsequent publication.
@@ 6 Conclusions and Future Work
In this work, we have presented an online framework which is flexible to varying demands from a user, and which is moderately fast even under comparatively adversarial circumstances. It is highly modular, allowing easy implementation of data extraction and analysis code for additional diagnostic devices. Though the code’s performance during experimental runtime was largely successful, we would like to also highlight the framework’s weaknesses, and how it could be improved and modified in the future.
@@ 6.1 Introducing Caching
Improvements could be made to execution speed by locally caching data from shots of interest. Often, it was desirable to plot data from the same shot multiple times in order to, for example, crop to regions of interest or adjust image contrast for visibility. This required multiple code runs, with each run limited by the speed at which data could be acquired over a network connection. In future, it would be useful to introduce an optional “caching” feature, allowing users to store data from a desired shot locally over a short period, with the option of “clearing” the cache when this data was no longer required.
@@ 6.2 Tagging by Cyclestamp
Experimental collaborations using the framework should aim to have some universally-recognized system of logging data by acquisition time. At CERN, as discussed in Section 3, a strong candidate would be the UNIX timestamp corresponding to each accelerator cycle (“cyclestamp”), as data filenames can be “stamped” by unique UNIX timestamp corresponding to beam extraction times. The framework codebase is easily adjusted to allow indexing by UNIX timestamp so long as it knows the position of the timestamp in the filename string. More sophisticated standards of synchronization are possible and preferable to the kind employed in this framework. For example, the White Rabbit project at CERN provides sub-nanosecond synchronization precision using the IEEE 1588-2008 Precision Time Protocol (PTP) [whiterabbit6070148]. Future experiments may wish to distinguish between multiple data acquisition events for the same beamline extraction if devices are expected to trigger more than once at the sub-second level.
However, in the construction and execution of our framework, bottlenecks in speed and performance emerged primarily due to delays in the process of saving files from the multitude of diagnostic devices (described in Section 2) to our cloud-based CERNBox/EOS storage system, and this delay was folded into our latency budget constraints. Our framework has no method for increasing the speed of caching diagnostic data to the CERNBox/EOS repository after beam extraction, which is a hardware performance concern rather than synchronization rate limitation. Modifying our framework to include sub-nanosecond synchronization schemes will fail to address this shortcoming. Indeed, during our experiment, timestamp resolution needed only to be lower than the interval between beamline extractions (see Section 3.1.1). Synchronization methods with precision on or below the order of are readily achievable using standard operating system methods for synchronization between devices connected over Wi-Fi, such as the Windows Time (W32Time) Service [MS_SupportBoundaryHighAccuracy], which employs NTP and can achieve synchronization if network latency does not exceed between devices.
Further improvements to the framework should focus on increasing the understanding of the underlying hardware, particularly the precise rate at which each diagnostic device logs data to the CERNBox/EOS repository, and the latency between beam extraction and data caching which subsequently emerges for each device. Revising the framework to give increased understanding and control over the underlying diagnostic hardware is utterly crucial in improving the system we have described in this technical note. Once again, these concerns cannot be addressed by simply improving the precision of the underlying synchronization protocol.
@@ 6.3 Avoiding Intermediate Cloud-Based Services
Though the CERN SWAN service allowed direct user access to the CERNbox service, it introduced several obstacles. SWAN has no native functionality for editing JSON file types. Furthermore, the Jupyter kernel introduces further overhead, and online code refactors in the framework modules are not recognized unless the kernel is restarted. Additionally, SWAN requires a stable connection to the CERN network, which is in-demand during working hours, introducing unpredictable latency. A possible future workaround for experiments could be saving data to a local computer, rather than cloud-based service, on which the framework has been installed.